Question 1 (uniform circular motion) A 1500 kg car is taking a left turn at a constant speed of 10 m/s. The radius of the turn is 12 meters. a. What is the centripetal acceleration on the car? b. How much force is require to hold the car in that turn? c. What force(s) is responsible for the centripetal force on the car?
Answer #1

Centripital acceleration a = v^2/r = 10×10÷12

= 8.33m/s2

Force required to hold the car is F = Ma = 1500×8.33 F =12500 N

Friction force is responsible for this centripital force. If there is no friction then car will start slipping on reaching the curve.

One other force that may be responsible for this centripital force component of gravitational force acting toward center of the curvature only if there is banking of road. If no banking then this force cant come into effect